In a 9 inch square cake pan and a 10 inch square cake pan, what is the difference in volume each pan will hold? Assume each pan is 3 inches high. Show step by step work. Include correct units with your solution.

Answer by checkley77(12844)

9*9*3=81*3=243 IN^3.
10*10*3=100*3=300 IN^3.
300-234=66 IN^3 IS THE DIFFERENCE IN VOLUME.